2e^-x + e^y = 3^x-y at 0,0


1
Expert's answer
2021-04-15T07:34:01-0400


Given,

"2e^{-x} + e^y = 3^{x-y}"


Differentiate with respect to x-


"-2e^{-x}+e^y\dfrac{dy}{dx}=3^{x-y}[1-\dfrac{dy}{dx}]"


then Rearranging the equation and we get-


"\dfrac{dy}{dx}=\dfrac{2e^{-x}+3^{x-y}}{e^y+3^{x-y}}"


"\dfrac{dy}{dx}_{(0,0)}=\dfrac{2e^0+3^0}{e^0+3^0}=\dfrac{2+1}{1+1}=\dfrac{3}{2}"